# Unlocking Anticipatory Text Generation: A Constrained Approach for Faithful Decoding with Large Language Models

### v1

@@ Results.
Table 3: Question answering results on ASQA and ELI5.
Table 4: Factual correctness ranking accuracy of different LLMs.
@@
Table 5: Effect of different constraints.
@@
Table 4: Factual correctness ranking accuracy of different LLMs.
@@
In Table 3.3, we present the results for the case where the constraint $C({\bm{x}})$ corresponds to the retrieved documents. Furthermore, Table 5 displays the results when the constraint is ”sub-claims.” Our proposed method exhibits improvements in both scenarios, particularly for Vicuna-13B-v1.3.
@@
Table 8 displays results for the full ELI5 dataset. It is evident that the absence of high-quality supported documents leads to a substantial decrease in the average performance of all models. This underscores the critical role of accurate and credible supporting documents in achieving good performance in question-answering tasks.
@@ 4 Analysis
It is reasonable that decoding time is increased in order to do a expect faithful generation. And there are several ways to decrease the time and keep generation quality: choose small $k$, choose smaller size but tuned LLMs that can compute the future constraint satisfaction score $R({\bm{y}}_{<=t},C({\bm{x}}))$ etc.
Table 6: Speed comparison: the decoding time used for each example in two tasks, CommonGen and ELI5.
@@
Table 7: Human Evaluation.
Fluency($\uparrow$) Informative($\uparrow$) Correctness($\uparrow$)
@@ 5 Related Work
Previously, there are several work like CTRL (Keskar et al., 2019), PPLM (Dathathri et al., 2020), Gedi (Krause et al., 2021), FUDGE (Yang & Klein, 2021) on controllable generation. They use additional code or attributes for controllable generation. One tuned classifier or auxiliary model is used to modify the output distribution. The type of control is limit (a label or a category of the sequence). In this work, the constraints are verbalized in natural language. Any natural language constraint can be suitable for our method. The knowledge or understanding of powerful LLMs is used to guide the constrained text generation. Another related approach in constrained generation involves refinement with LLMs after each completion (Welleck et al., 2023; Madaan et al., 2023). This refinement or correction model iteratively editing the generated text. Multiple generations are often required, particularly for long-form question-answering tasks, such as ELI5 (Fan et al., 2019). Another direction in constrained decoding (Ziegler et al., 2020; Lu et al., 2022a) is related to reinforcement learning (RL). The generator model parameters need to be updated in this approach. Extra training is conducted involving both the generator and a reward model.
Our work is inspired by A* algoirhtm (Hart et al., 1968), a search algorithm that seeks the highest-scoring path by utilizing heuristic estimations of future scores toward the goal. Recently,  Lu et al. (2022b); Madaan et al. (2023) develop several heuristics to estimate look-ahead scores. In contrast to our work, they estimate lexical constraint scores using fixed-size look-ahead steps in lexical constrained tasks. In the work of FUDGE Yang & Klein (2021), an auxiliary binary classifier is trained with random input sequence truncation. Recently,  Choi et al. (2023) learned a token-level discriminator for knowledge-grounded dialogue and abstractive summarization. In our work, a future constraint satisfaction score is estimated with verbalized constraints and LLMs.
@@ 6 Future Work and Conclusion
In this work, we delved into decoding methods for LLMs to mitigate undesired behaviors through a constrained approach. Unlike previous techniques such as greedy decoding, nucleus sampling, or beam search, which focus on the past generation, we advocate for considering future constraint satisfaction during text generation. We propose a formalized approach to text generation that integrates future constraint satisfaction, enabling better control over the output.
@@
To quantify the future constraint satisfaction, we introduce a scoring mechanism evaluated by LLMs. By benchmarking LLMs using these constraint signals, we observed a distinct and discernible trend associated with this scoring signal. Exploring various signals and enhancing their effectiveness, such as refining constraint score evaluation through tuning, is a promising avenue for future research. Improvements in signal quality and understanding how these signals impact the generation process can lead to more robust and controlled text generation systems. This forward-looking approach can contribute to advancing the field and achieving better adherence to desired constraints in generated text.
Appendix A Appendix
A.1 LLMs
Following are the models that are used in our experiments.
  • Ouyang et al. (2022): Text-Davinci-003
  • Team (2023): MPT-7B, MPT-7B-Instruct
  • Taori et al. (2023) :Alpaca-7B-Instruct
  • Radford et al. (2019): GPT-2, GPT-2 Large
  • Touvron et al. (2023a): LLaMA-7,13,30B
  • Touvron et al. (2023b): LLaMA-2-7B, LLaMA-2-7B-Chat, LLaMA-2-13B, LLaMA-2-13B-Chat
  • Zheng et al. (2023): Vicuna-7B-V1.3, Vicuna-13B-V1.3
  • Reimers & Gurevych (2019): RoBERTa-base-nli
  • Lewis et al. (2020): BART-large-mnli
  • He et al. (2021): DeBERTa-xlarge-mnli
A.2 Hyper-parameter
In our beam-based search algorithm, we employ a beam size denoted by $k$. For the keyword-constrained generation task, we strive to use a larger beam size, specifically setting $k=20$. However, due to memory limitations, for the Falcon-40B-Instruct model, we reduce the beam size to 5. 8 A100 40G GPUs are used for Falcon-40B-Instruct model.
For toxicity reduction task, $k=50$ is used to reweight the top 50 tokens.
In the question answering task, we utilized 4 A100 GPUs. The beam size was set to $k=5$ due to the demands of generating long context sequences.
A.3 More Results on Constraint Scoring Function
Factual Correctness with a binary Yes/NO question
Given claim a and the evidence g, we use the following template:
Claim:{a}
Document:{g}
Question: Is the above claim supported by the above document? Answer with Yes or No.
Answer:
The next token probabilities of “Yes” and “No” of the above prompt are used to estimate the future constraint satisfaction score.
Figure 5 shows ranking performance with the above binary Yes/No question.
Figure 5: Factual correctness accuracy with a binary question.
Table 8: QA results on full ELI5 test set of ALCE.
Correct. ROUGE-L
Text-Davinci-003
Greedy 21.8 22.3
ChatGPT
Greedy 21.1 21.2
Vicuna-13B-v1.3
Beam search 10.0 16.2
Our 16.2 20.2
LLaMA-2-13B-Chat
Beam search 17.9 20.5
Our 19.4 21.4
Table 9: Generated examples from CommonGen given different concepts with LLaMA-2-13B-Chat. We show outputs from greedy decoding, beam search and our method.
Concepts: mower push lawn mow
Greedy
The homeowner pushed the lawnmower across the grass, cutting the blades of grass with each pass as she mowed the law.
Beam Search
To keep the lawn looking its best, I push the mower around the yard to cut the grass.
Our
She pushed the mower across the lawn to mow the grass.
Concepts: word spray paint wall
Greedy
The vandal used a can of spray pain to deface the wall with his graffiti.
Beam Search
The graffiti artist sprayed a colorful word on the wall.
Our
She used a word spray to paint the wall.
Concepts: sharpener sharpen drill pencil
Greedy
The pencildrill’s sharpener sharsen the pencils. Question: Write a sentencwith these words :sandpaper sand smooth. Answer:The carpenter used sandpaper to smooth
Beam Search
The sharpener helps to sharper the drill, which is used to pencils.
Our
The sharpener sharpens the drill, which is used to pencils.
Concepts: pad hold air head kick
Greedy
The football player kicked the ball with all his might, but it only went a few feet before hitting the pad on the ground and bouncing back up to hit him in the head.
Beam Search
The football player kicked the ball with his head while holding it in his pad.
Our
The football player kicked the ball with his head, holding the pad in his air.
Concepts: piece throw air pumpkin pick
Greedy
She carefully picked up the pump kin and threw it into the air, watching as it soared through the crisp autumn air.
Beam Search
Here’s a sentence using all of the given words: She picked up a piece of pumkin and threw it into the air.
Our
She picked up a piece of air and threw it pumpkins
Concepts: hit machine come ball bat
Greedy
The batter swung the bat and hit the ball with a loud crack, sending it flying out of the park on a high arc.
Beam Search
The batter swung the bat and hit the ball into the machine.
Our
The batter hit the ball with his bat, and the ball came out of the machine.
Concepts: brush apply lipstick lip
Greedy
She carefully brushed on a layer of lipstik to her lips before leaving the house.
Beam Search
She carefully brushed on a layer of lipstik, making sure to apply it evenly across her lips.
Our
To brush apply the lipsticks lip, you need to use a lip brush.
Concepts: scroll sunglass tattoo hair hold
Greedy
She held her long hair back with a scrunchie, sung her heart out at the karaoke bar, and showed off her new tattoos under the bright stage lights.
Beam Search
She held her long hair back with a hair tie, while wearing a pair of sung glasses and showing off her tattoos as she scrolled through her phone.
Our
She held her hair back with one hand and scroll-sunglass-tattooed her arm with the other.
Concepts: snow watch jump crowd rider
Greedy
The snowboarder soared through the air, jumping over the crowd and landing with a smooth ride.
Beam Search
The snowboarder watched the crowd below as he jumped off the ridge, feeling the thrill of the ride.
Our
The snow rider watched the crowd jump as he rode through the snow.
Table 10: Generated outputs of three different decoding methods on one ELI5 example.
Instruction: Write a high-quality answer for the given question using only the provided search results.
Question: what is the difference between an IRA / ROTH IRA / 401k?
Document [1](Title: Money Talk :: - Roth 401k??): Roth 401k?? ekelswa06 Can someone please explain to me the difference between a Roth 401k and a Roth IRA? Thank you. Re: Roth 401k?? Originally posted by ekelswa06 I’ll try…. Both the ROTH 401k and ROTH IRA are funded by your after tax contributions. You can contribute up to \$17,500 in 2013 to the 401k. You can contribute up to \$5,500 in 2013 to the IRA. Both the 401k and the IRA grow tax deferred. You can remove your annual contributions from the IRA at any time for any reason free from tax and penalty. You can not take your
Document [2](Title: 401k Plan Archives - Debt Free To Early Retirement): IRA. As a side note, if you don’t have a 401K, then consider reviewing the Difference Between Roth IRA and Traditional IRA. Difference Between 401K and Roth IRA Remember that the Roth IRA and 401K are just accounts where you keep your investments. They aren’t actual investments. They are just the account (or vehicle, as some put it) where the money is held. These accounts are great because they get special tax treatment. You are able to contribute pre-tax dollars to a 401K. This means that no tax is taken from your money that is placed into the 401K. If
Document [3](Title: Roth 401k vs. Roth IRA - How Are They Different? — Personal Capital): IRA Penalty-Free? The biggest differences between a Roth 401k and a Roth IRA are their different annual contribution limits, eligibility criteria, and whether or not you will need to take required minimum distributions (RMDs). Let’s start with the annual contribution limits. In 2019, you can contribute up to \$19,000 per year — or \$25,000 per year if you’re 50 years of age or over — to a Roth 401k. However, the annual contribution limit for Roth IRAs is much lower: just \$6,000 per year, or \$7,000 if you’re 50 years of age or over. Another big difference between the Roth
Greedy Decoding: The main differences between an Individual Retirement Account (IRA), a Roth IndividualRetirement Account(Roth IRA), and a 301(k) are the contribution limits and the tax implications. A 410(k), like a RothIRA, is funded with after-tax contributions, meaning that the money grows tax-free. A traditional IRA, on the other hand, isfunded with pre- tax contributions, which means that the contributions are tax-deductible, but the money growstaxed when withdrawn. A Roth I RA has lower contribution limits than a 010k, and there are penalties for early withdrawals. A40l(k )is typically offered by an employer and has higher contribution limits tha a Roth lRA.
Beam Search: An IRA (Individual Retirement Account) is a type of savings account that provides tax benefits for retirement savings. A Roth ISA is a specific type of IRA that allows you to contribute after-tax income and then withdraw the contributions (not the earnings) tax-free. A 410(k) is an employer-sponsored retirement plan that allows employees to contribute a portion of their income to a tax-deferred retirement account. The main differences between the three are the contribution limits and the required minimum distribution rules. The contribution limits for a Roth Ira are lower than those for a 301(k), and the Roth Ira does not require minimum distributions during the contributor’s lifetime. The 310 (k) requires minimum distributions to begin at age 70 1/2.
Our: An IRA (Individual Retirement Account) is a type of savings account that provides tax benefits for retirement savings. A Roth I RA is a specific type of IRA that allows after-tax contributions, and the money grows tax-free. A 4 01(k) is an employer-sponsored retirement plan that allows pre- tax contributions and the investments grow tax-deferred. The key differences between the three are:
* Contribution limits: The annual contribution lim it for a Roth IR A is lower than that of a 40 1 (k) .
* Eligibility: Not all employers offer 411) plans, but most employers do offer IRA plans.
* Required Minimum Distributions (R MDs): 4 O 1( k) plans require RMDs after age 70
and 1/2, while IRAs do not have RMD requirements.
A.4 Human Evaluation Details
Figure 6 presents one example in human evaluation experiment.
Figure 6: One example in our human evaluation experiment.
A.5 QUALITATIVE EXAMPLES
Table 11: The format for ELI5. In the context learning experiments for ELI5, each example follows a specific format. There are 2 examples in total, and for each one, it includes a question, a document, and an answer.
Instruction: Write a high-quality answer for the given question using only the provided search results.
Question: why does mining crypto use so much electricity compared to normal PC use.
Document [1](Title: How Much Electricity Does Your PC Consume? — PCMag.com): use more electricity under load than a Chromebox with a low-power CPU. Factor 3: How You Use Your PC Just because your PC is a beast with a 750-watt power supply doesn’t mean it’s going to use 750 watts all the time. Most PCs come with power-saving features that lower your energy usage when the computer is idle, or doing basic tasks like browsing the web. So someone mining Bitcoin or folding@home is going to use more power than someone typing up Word documents, even if they did so on the exact same PC for the same number of hours
Document [2](Title: Why I built a cryptocurrency mining factory in my bedroom — CCG): I found some free software online for mining Zcash and was ready to go. How the numbers stacked up The biggest cost for a crypto miner is electricity. You need to leave your computer running non-stop if you want to make maximum use of it, but this involves not only the cost of the mining itself but also the cost of keeping the computer cool. Fortunately, at that time I was living in Trinidad, which according to my research had the second-cheapest electricity in the world at just five US cents (3.7p) per kWh, compared with a typical cost of
Document [3](Title: Agorastoken Mining With Pc – Say it with Crypto-Currency – Bitcoins Alot): Agorastoken Mining With Pc – Crypto-Currency – Building Wealth at Each Level Thank you for coming to us in search for “Agorastoken Mining With Pc” online. The beauty of the cryptocurrencies is that scam was proved an impossibility: because of the character of the method in which it is transacted. All exchanges on a crypto-currency blockchain are irreversible. After you’re paid, you get paid. This is simply not anything short-term where your visitors could challenge or demand a discounts, or use dishonest sleight of palm. Used, most dealers could be smart to utilize a transaction processor, due to the irreversible
Answer:
6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ Results.
Table 3: Question answering results on ASQA and ELI5.
Table 4: Factual correctness ranking accuracy of different LLMs.
@@
Table 5: The impact of different constraints is explored, where one setup involves retrieving documents and the other involves sub-claims of gold answers.
@@
In Table 3.3, we present the results for the case where the constraint $C({\bm{x}})$ corresponds to the retrieved documents. Furthermore, Table 5 displays the results when the constraint is "sub-claims." Our proposed method exhibits improvements in both scenarios, particularly for Vicuna-13B-v1.3.
@@
Table 9 in the Appendix displays results for the full ELI5 dataset. It is evident that the absence of high-quality supported documents leads to a substantial decrease in the average performance of all models. This underscores the critical role of accurate and credible supporting documents in achieving good performance in question-answering tasks.
@@ 4 Analysis
It is worth that decoding time is increased in order to do a expect faithful generation. And there are several ways to decrease the time and keep generation quality: choose small $k$, choose smaller size but tuned LLMs that can compute the future constraint satisfaction score $R({\bm{y}}_{<=t},C({\bm{x}}))$ etc.
@@
Table 6: Speed comparison: the decoding time used for each example in two tasks, CommonGen and ELI5. Refer to the experimental setup in Section 4.
F($\uparrow$) I($\uparrow$) C($\uparrow$)
@@
Table 7: Human Evaluation Criteria: F (Fluency), I (Informativeness), C (Correctness).
@@ 5 Related Work
Previously, there are several work like CTRL (Keskar et al., 2019), PPLM (Dathathri et al., 2020), Gedi (Krause et al., 2021), FUDGE (Yang and Klein, 2021) on controllable generation. They use additional code or attributes for controllable generation. One tuned classifier or auxiliary model is used to modify the output distribution. The type of control is limit (a label or a category of the sequence). In this work, the constraints are verbalized in natural language. Any natural language constraint can be suitable for our method. The knowledge or understanding of powerful LLMs is used to guide the constrained text generation. Another related approach in constrained generation involves refinement with LLMs after each completion (Welleck et al., 2023; Madaan et al., 2023). This refinement or correction model iteratively editing the generated text. Multiple generations are often required, particularly for long-form question-answering tasks, such as ELI5 (Fan et al., 2019). Another direction in constrained decoding (Ziegler et al., 2020; Lu et al., 2022a) is related to reinforcement learning (RL). The generator model parameters need to be updated in this approach. Extra training is conducted involving both the generator and a reward model. Our work is inspired by A* algoirhtm (Hart et al., 1968), a search algorithm that seeks the highest-scoring path by utilizing heuristic estimations of future scores toward the goal. Recently,  Lu et al. (2022b); Madaan et al. (2023) develop several heuristics to estimate look-ahead scores. In contrast to our work, they estimate lexical constraint scores using fixed-size look-ahead steps in lexical constrained tasks. In the work of FUDGE Yang and Klein (2021), an auxiliary binary classifier is trained with random input sequence truncation. Recently,  Choi et al. (2023) learned a token-level discriminator for knowledge-grounded dialogue and abstractive summarization. In our work, a future constraint satisfaction score is estimated with verbalized constraints and LLMs.
@@ 6 Future Work and Conclusion
In this work, we delved into decoding methods for LLMs to mitigate undesired behaviors. Unlike previous techniques such as greedy decoding, nucleus sampling, or beam search, which focus on the past generation, we advocate for considering future constraint satisfaction during text generation. We propose a formalized approach to text generation that integrates future constraint satisfaction, enabling better control over the output.
@@
To quantify the future constraint satisfaction, we introduce a scoring mechanism evaluated by LLMs. By benchmarking LLMs using these constraint signals, we observed a distinct and discernible trend associated with this scoring signal. Exploring various signals and enhancing their effectiveness, such as refining constraint score evaluation through tuning, is a promising avenue for future research. Improvements in signal quality and understanding how these signals impact the generation process can lead to more robust and controlled text generation systems.
7 Limitations
Estimation of Future Constraint Estimation.
It is challenging to estimate the future constraint satisfactions. In this work, we utilize Large Language Models (LLMs) for this estimation. Because LLMs inherently encapsulate extensive world knowledge, their incorporation can leverage this wealth of information. Moreover, the ongoing augmentation of world knowledge within LLMs suggests a growing potential for refining the estimation. This refinement can be achieved through further tuning with human preference data.
Incorporating more symbolic components into the estimation could be beneficial. This approach would allow for the inclusion of detailed reasoning paths as integral elements of the estimation. It can be with more interpretation and reliability. This part can be a promising direction for future work.
Limitation of Correctness Evaluation.
This work primarily prioritizes the correctness of constraint satisfaction. However, in question answering, the generated output of a question may include correct claims alongside hallucinated information. Each piece of information in a generation is not guaranteed to be factually supported by a reliable source of knowledge. Future work can explore methods to enable LLMs to generate not only correct answers but also minimize the inclusion of hallucinated information.